OpenStudy (anonymous):

factor 49n^2 + 56n + 7

OpenStudy (kittiwitti1):

\[49n^{2}+56n+7\]\[\text{find the greatest common factor of 49, 56 and 7}\]\[\text{so basically what single number can be multiplied}\]\[\text{by a certain other number to get each number?}\]

OpenStudy (anonymous):

7

OpenStudy (kittiwitti1):

do you know what to do now then?

OpenStudy (anonymous):

nope

OpenStudy (kittiwitti1):

how do you factor something mero?

OpenStudy (anonymous):

(7n + something) (n +something)

OpenStudy (kittiwitti1):

well there is an easier way but that works too i guess\[\text{GCF}(\frac{49}{\text{GCF}}n^{2}+(\frac{56}{\text{GCF}})n+(\frac{7}{\text{GCF}})\]GCF means greatest common factor

OpenStudy (kittiwitti1):

then take out the stuff outside parentheses and factor the stuff inside. the smaller the numbers the easier to solve mero~ o vo

OpenStudy (anonymous):

how would you do that for this

OpenStudy (kittiwitti1):

i just explained that to you

OpenStudy (anonymous):

no i know but for this example

OpenStudy (kittiwitti1):

ok. what's the GCF ?

OpenStudy (anonymous):

is it 7

OpenStudy (kittiwitti1):

so replace "GCF" with 7 in the equation i gave you then take out the 7 outside the parentheses and keep the equation inside the parentheses

OpenStudy (anonymous):

7n^2 + 8n + 1

OpenStudy (kittiwitti1):

yes, \[7n^{2}+8n+1\]

OpenStudy (kittiwitti1):

can you solve now?

OpenStudy (anonymous):

(7n + 1 ) (n + 1)

OpenStudy (kittiwitti1):

i'm kinda rusty with the actual factoring stuff when n^2 has a coefficient sorry

OpenStudy (kittiwitti1):

I got this:\[(7n+7)(n+1)\]